# Dynamic Tuning Towards Parameter and Inference Efficiency for ViT Adaptation

### v1

@@ 4.3 VTAB-1K Results
To evaluate the adaptation performance when the training data is limited, we adapt the VTAB-1K [74] benchmark. Following exiting works, we reduce the bottleneck dimension $d$ to 8, resulting in only 0.16M extra parameters. We vary the activation rate $r$ in DyT in the range [0.5, 0.7, 0.9] and conduct fine-tuning, obtaining three models with different inference costs.
The results are presented in Table 5. Although previous methods, such as ConvPass [36] and Res-Tuning [35], achieve satisfactory performance, they do not improve the efficiency during inference and even introduce additional FLOPs compared with full fine-tuning. In contrast, with only 12.54 GFLOPs, about 71% of the cost in original ViT-B, our method has outperformed most previous methods e.g. LoRA [32] and VPT [34], by a large margin. Furthermore, as FLOPs increase, DyT continues to exhibit superior performance than other methods. These experimental results validate that DyT effectively improves parameter efficiency and inference efficiency while maintaining the adaptation performance.
Table 5: Performance and efficiency comparison on VTAB-1K. Our approach achieves significant adaptation performance while simultaneously enhancing both parameter and inference efficiency. “Group Mean” indicates the averaged accuracy of three groups. “Full tuning” indicates fine-tuning all parameters. “Frozen” means that all parameters in the backbone keep frozen. “Param. (M)” denotes the number of trainable parameters in backbones. “FLOPSs (G)” is the average FLOPs across all datasets. Bold font and underline denote the best and the second-best performance respectively.
Natural Specialized Structured Structured
@@
Full tuning 68.9 87.7 64.3 97.2 86.9 87.4 38.8 79.7 95.7 84.2 73.9 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1 68.96 85.8 17.58
Frozen 63.4 85.0 63.2 97.0 86.3 36.6 51.0 78.5 87.5 68.6 74.0 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2 57.64 0 17.58
@@
Adapter [31] 69.2 90.1 68.0 98.8 89.9 82.8 54.3 84.0 94.9 81.9 75.5 80.9 65.3 48.6 78.3 74.8 48.5 29.9 41.6 73.85 0.16 17.61
BitFit [72] 72.8 87.0 59.2 97.5 85.3 59.9 51.4 78.7 91.6 72.9 69.8 61.5 55.6 32.4 55.9 66.6 40.0 15.7 25.1 65.21 0.10 17.58
LoRA [32] 67.1 91.4 69.4 98.8 90.4 85.3 54.0 84.9 95.3 84.4 73.6 82.9 69.2 49.8 78.5 75.7 47.1 31.0 44.0 74.60 0.29 17.58
VPT [34] 78.8 90.8 65.8 98.0 88.3 78.1 49.6 81.8 96.1 83.4 68.4 68.5 60.0 46.5 72.8 73.6 47.9 32.9 37.8 71.96 0.53 18.30
SSF [36] 69.0 92.6 75.1 99.4 91.8 90.2 52.9 87.4 95.9 87.4 75.5 75.9 62.3 53.3 80.6 77.3 54.9 29.5 37.9 75.69 0.20 17.58
NOAH [76] 69.6 92.7 70.2 99.1 90.4 86.1 53.7 84.4 95.4 83.9 75.8 82.8 68.9 49.9 81.7 81.8 48.3 32.8 44.2 75.48 0.36 17.58*
ConvPass [36] 72.3 91.2 72.2 99.2 90.9 91.3 54.9 84.2 96.1 85.3 75.6 82.3 67.9 51.3 80.0 85.9 53.1 36.4 44.4 76.56 0.33 17.64
AdaptFormer [10] 70.8 91.2 70.5 99.1 90.9 86.6 54.8 83.0 95.8 84.4 76.3 81.9 64.3 49.3 80.3 76.3 45.7 31.7 41.1 74.75 0.16 17.61
FacT-TT [37] 71.3 89.6 70.7 98.9 91.0 87.8 54.6 85.2 95.5 83.4 75.7 82.0 69.0 49.8 80.0 79.2 48.4 34.2 41.4 75.30 0.04 17.58
Res-Tuning [35] 75.2 92.7 71.9 99.3 91.9 86.7 58.5 86.7 95.6 85.0 74.6 80.2 63.6 50.6 80.2 85.4 55.7 31.9 42.0 76.32 0.51 17.67
@@
4.4 Effectiveness on Image Datasets with Sufficient Training Data
Although the results from VTAB-1K have proven the superiority of our approach, we extend our investigation to complete image datasets, to evaluate the adaptation performance with sufficient training data. The results are demonstrated in Table 6. We further explore a straightforward approach, “Dynamic full tuning”, which has a token dispatcher as the same in DyT and is fine-tuned with all parameters. We observe that its performance becomes unstable and drops significantly on some datasets e.g. CIFAR-100 and Food101. This phenomenon may arise due to the potential adverse impact of dynamic dispatch on the pre-trained parameters during full-tuning adaptation, thereby validating the importance of DyT. In this data-sufficient scenario, although our method achieves performance slightly below that of full tuning and AdaptFormer, it brings a significant reduction in FLOPs.
4.5 Scaling Token Counts from Images to Videos
@@
4.6 Generalization on Semantic Segmentation
@@
Table 6: Results on three tasks. We present top-1 accuracy for both image and video datasets, while reporting mIOU scores for semantic segmentation datasets ADE20K and COCO-stuff. “Avg.” denotes the average results from the corresponding two datasets. The FLOPs are evaluated on CIFAR-100, K400 and ADE20K.
  Method Params. $\downarrow$ Image Datasets Video Datasets Semantic Segmentation Datasets
FLOPs $\downarrow$ Image Accuracy (%) FLOPs $\downarrow$ Video Accuracy (%) FLOPs $\downarrow$ mIOU
(M) (G) CIFAR-100 SVHN Food101 Avg. (G) K400 SSv2 Avg. (G) ADE20K COCO-stuff Avg.
@@
Full tuning 85.80 17.58 90.91 97.29 90.69 92.96 142.53 75.48 45.22 60.35 605.36 47.66 45.89 46.78
Linear 0 17.58 85.87 56.29 88.07 76.74 142.53 69.04 27.64 48.34 605.36 43.86 44.46 44.16
Dynamic-Full 85.80 12.24 83.43 96.90 84.46 88.26 107.67 74.63 44.94 59.79 582.46 45.74 45.20 45.47
@@
AdaptFormer [10] 1.19 17.81 92.03 97.23 90.84 93.36 144.39 75.53 45.36 60.45 606.57 47.52 45.33 46.43
LoRA [32] 1.19 17.58 91.42 97.36 90.48 93.08 142.53 75.48 45.62 60.55 605.36 47.34 45.53 46.44
VPT [34] 0.07 18.32 91.64 95.72 90.41 92.59 148.44 73.46 38.17 55.82 606.36 46.38 44.87 45.63
@@ 5 Discussion and Conclusion
Limitations and future works. DyT is currently designed for vision tasks. We believe it would be interesting to combine vision backbones with large language models e.g. [62] to build efficient large multi-modal models e.g. [45, 80]. DyT could be further developed to be compatible with multi-modal models, which would be our future direction.
Appendix 0.A Appendix
0.A.1 Details about Gumbel-Sigmoid

### v2

@@ 4.3 VTAB-1K Results
Comparison with PEFT methods.
To evaluate the adaptation performance when the training data is limited, we adapt the VTAB-1K zhai2019large benchmark, which a widely employed to evaluate the performance of adaptation tasks. Following exiting works, we reduce the bottleneck dimension $d$ to 8, resulting in only 0.16M extra parameters. We vary the activation rate $r$ in DyT in the range [0.5, 0.7, 0.9] and conduct fine-tuning, obtaining three models with different inference costs.
The results are presented in Table 3. Although previous methods, such as ConvPass jie2022convolutional and Res-Tuning jiang2023res , achieve satisfactory performance, they do not improve the efficiency during inference and even introduce additional FLOPs compared with full fine-tuning. In contrast, with only 12.54 GFLOPs, about 71% of the cost in original ViT-B, our method has outperformed previous methods e.g. LoRA hu2021lora and VPT jia2022visual , by a large margin. Remarkably, the DyT model with $r=0.9$ does not surpass the performance of the DyT model with $r=0.7$. This observation suggests the presence of computational redundancy in the transformer after adaptation, which further validates the significance of dynmaic tuning. These experimental results validate that DyT effectively improves parameter efficiency and inference efficiency while maintaining the adaptation performance.
Table 3: Performance and efficiency comparison on VTAB-1K. “Group Mean” indicates the averaged accuracy of three groups. “Params. (M)” denotes the number of trainable parameters in backbones. “FLOPSs (G)” is the average FLOPs across all datasets. Bold font and underline denote the best and the second-best performance respectively.
Natural Specialized Structured
@@
Full tuning 68.9 87.7 64.3 97.2 86.9 87.4 38.8 79.7 95.7 84.2 73.9 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1 68.96 85.80 17.58
Frozen 63.4 85.0 63.2 97.0 86.3 36.6 51.0 78.5 87.5 68.6 74.0 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2 57.64 0.00 17.58
@@
Adapter houlsby2019parameter 69.2 90.1 68.0 98.8 89.9 82.8 54.3 84.0 94.9 81.9 75.5 80.9 65.3 48.6 78.3 74.8 48.5 29.9 41.6 73.85 0.16 17.61
BitFit zaken2021bitfit 72.8 87.0 59.2 97.5 85.3 59.9 51.4 78.7 91.6 72.9 69.8 61.5 55.6 32.4 55.9 66.6 40.0 15.7 25.1 65.21 0.10 17.58
LoRA hu2021lora 67.1 91.4 69.4 98.8 90.4 85.3 54.0 84.9 95.3 84.4 73.6 82.9 69.2 49.8 78.5 75.7 47.1 31.0 44.0 74.60 0.29 17.58
VPT jia2022visual 78.8 90.8 65.8 98.0 88.3 78.1 49.6 81.8 96.1 83.4 68.4 68.5 60.0 46.5 72.8 73.6 47.9 32.9 37.8 71.96 0.53 18.30
SSF jie2022convolutional 69.0 92.6 75.1 99.4 91.8 90.2 52.9 87.4 95.9 87.4 75.5 75.9 62.3 53.3 80.6 77.3 54.9 29.5 37.9 75.69 0.20 17.58
NOAH zhang2022neural 69.6 92.7 70.2 99.1 90.4 86.1 53.7 84.4 95.4 83.9 75.8 82.8 68.9 49.9 81.7 81.8 48.3 32.8 44.2 75.48 0.36 17.58*
ConvPass jie2022convolutional 72.3 91.2 72.2 99.2 90.9 91.3 54.9 84.2 96.1 85.3 75.6 82.3 67.9 51.3 80.0 85.9 53.1 36.4 44.4 76.56 0.33 17.64
AdaptFormer chen2022adaptformer 70.8 91.2 70.5 99.1 90.9 86.6 54.8 83.0 95.8 84.4 76.3 81.9 64.3 49.3 80.3 76.3 45.7 31.7 41.1 74.75 0.16 17.61
FacT-TT jie2023fact 71.3 89.6 70.7 98.9 91.0 87.8 54.6 85.2 95.5 83.4 75.7 82.0 69.0 49.8 80.0 79.2 48.4 34.2 41.4 75.30 0.04 17.58
Res-Tuning jiang2023res 75.2 92.7 71.9 99.3 91.9 86.7 58.5 86.7 95.6 85.0 74.6 80.2 63.6 50.6 80.2 85.4 55.7 31.9 42.0 76.32 0.51 17.67
@@
Dynamic tuning achieves actual acceleration.
As a hardware-agnostic metric, FLOPs is not the most suitable choice to evaluate the inference efficiency across diverse platforms. The real inference speed has usually been ignored in previous PEFT methods. Here, we adopt two GPUs (Tesla V100 and Tesla T4) and a CPU Xeon(R) Platinum 8163 to comprehensively evaluate the efficiency of our methods and three representative PEFT methods, including LoRA hu2021lora , AdaptFormer chen2022adaptformer , and VPT jia2022visual . The batch size during inference is set to 512 and 32 for GPUs and the CPU, respectively. The results, as summarized in Table 4, reveal that our method achieves better performance while effectively accelerating inference speed compared to existing PEFT methods on different platforms.
Table 4: Comparison of throughput. “VTAB-1K Accuracy $\uparrow$” denotes the averaged accuracy of three dataset groups in VTAB-1K zhai2019large benchmark.
  Method VTAB-1K FLOPs (G) $\downarrow$ V100 T4 Xeon(R) 8163
Accuracy $\uparrow$ Throughput (img/s) $\uparrow$ Throughput (img/s)$\uparrow$ Throughput (img/s) $\uparrow$
Full tuning 68.96 17.58 806.24 435.41 2.12
LoRA hu2021lora 74.60 17.58 806.24 435.41 2.12
AdaptFormer chen2022adaptformer 74.75 17.61 767.30 400.42 1.97
VPT jia2022visual 71.96 18.30 762.55 392.13 1.95
DyT $r=0.5$ 77.14 12.54 912.30 524.93 3.89
Comparison and compatibility with methods for efficient transformers.
@@
Then, we explore the compatibility of our method with token pruning methods. Specifically, we combine DyT with ToMe bolya2022token , a training-free technique that progressively prunes tokens through token merging. From the results in Table 5, we find that ToMe can further enhance the throughput of DyT while maintaining accuracy. This proves the potential of our methods to be combined with existing token pruning methods e.g. bolya2022token ; chen2023diffrate ; liang2022not . Additionally, we apply ToMe to full tuning and AdaptFormer chen2022adaptformer in Table 3 and observe sub-optimal accuracy and throughput. These findings highlight that directly applying ToMe after fine-tuning or parameter-efficient fine-tuning is less effective compared to the proposed approach.
Table 5: Comparison with efficient transformers. The throughput is measured on a Tesla V100 GPU. “Params. (M) $\downarrow$ ” denotes the number of trainable parameters in backbones.
  Method VTAB-1K FLOPs (G) $\downarrow$ Param. (M) $\downarrow$ Throughput (img/s) $\uparrow$
Accuracy $\uparrow$
DynamicViT rao2021dynamicvit 60.10 14.05 88.70 1010.40
DynamicViT+AdaptFormerchen2022adaptformer 75.48 14.23 3.10 954.82
EViT liang2022not 67.42 11.62 85.80 1233.45
EViT+AdaptFormerchen2022adaptformer 74.63 11.77 0.16 1152.38
Full tuning + ToMe bolya2022token 68.68 13.12 85.80 989.29
AdaptFormer chen2022adaptformer + ToMe bolya2022token 74.30 13.29 0.16 941.70
DyT $r=0.5$ 77.14 12.54 0.16 912.39
DyT $r=0.5$ + ToMe bolya2022token 76.60 9.85 0.16 1114.70
4.4 Further Exploration
Effectiveness on image datasets with sufficient training data.
Although the results from VTAB-1K benchmark have proven the superiority of our approach, we extend our investigation to complete image datasets, to evaluate the adaptation performance with sufficient training data. We conduct experiments on 6 datasets including CIFAR-100 krizhevsky2009learning , SVHN goodfellow2013multi , Food-101 bossard2014food , Air maji2013fine , Pet parkhi2012cats , and Car gebru2017fine . The results are demonstrated in Table 6. We further explore a straightforward approach, “Dynamic-Full”, which has a token dispatcher as the same in DyT and is fine-tuned with all parameters. We observe that its performance becomes unstable and drops significantly on some datasets e.g. CIFAR-100 and Food-101. This phenomenon may arise due to the potential adverse impact of dynamic dispatch on the pre-trained parameters during full-tuning adaptation, thereby validating the importance of DyT. In this data-sufficient scenario, although our method achieves performance slightly below that of full tuning and AdaptFormer, it brings a significant reduction in FLOPs.
Scaling token counts from images to videos.
@@ 4.4 Further Exploration
Table 6: Results on image and video tasks. “Avg.” denotes the average results from the corresponding task. The FLOPs are evaluated on CIFAR-100 and K400.
  Method Params. $\downarrow$ Image Datasets Video Datasets
FLOPs $\downarrow$ Image Accuracy (%) FLOPs $\downarrow$ Video Accuracy (%)
(M) (G) CIFAR-100 SVHN Food-101 Air Pet Car Avg. (G) K400 SSv2 Avg.
@@
Full tuning 85.80 17.58 90.91 97.29 90.69 80.53 93.11 88.71 90.21 142.53 75.48 45.22 60.35
Linear 0 17.58 85.87 56.29 88.07 40.51 92.78 52.85 69.40 142.53 69.04 27.64 48.34
Dynamic-Full 85.80 12.24 83.43 96.90 84.46 62.50 75.16 70.48 78.82 107.67 74.63 44.94 59.79
@@
AdaptFormer chen2022adaptformer 1.19 17.81 92.03 97.23 90.84 78.23 94.46 87.33 90.02 144.39 75.53 45.36 60.45
LoRA hu2021lora 1.19 17.58 91.42 97.36 90.48 76.32 93.62 87.00 89.36 142.53 75.48 45.62 60.55
VPT jia2022visual 0.07 18.32 91.46 95.72 90.41 68.91 93.92 80.72 86.88 148.44 73.46 38.17 55.82
@@ 5 Discussion and Conclusion
DyT is currently designed for vision tasks. We believe it would be interesting to combine vision backbones with large language models e.g. touvron2023llama to build efficient large multi-modal models e.g. liu2024visual ; zhu2023minigpt . DyT could be further developed to be compatible with multi-modal models, which would be our future direction.
Acknowledgements
This work was supported by Damo Academy through Damo Academy Research Intern Program. Yang You’s research group is being sponsored by NUS startup grant (Presidential Young Professorship), Singapore MOE Tier-1 grant, ByteDance grant, ARCTIC grant, SMI grant (WBS number: A-8001104-00-00), Alibaba grant, and Google grant for TPU usage.
Appendix A Appendix
We organize our appendix as follows.
  • In Section A.1, we present frequently asked questions along with their corresponding answers.
  • In Section A.2, we detail the difference between our method and other previous works.
  • In Section A.3, we present more analysis on the proposed method.
  • In Section A.4, we report the results on semantic segmentation datasets.
  • In Section A.5, we report the performance on object detection and instance segmentation.
  • In Section A.6, we verify the effectiveness of complete model and distillation during adaption.
  • In Section A.7, we provide the details and a formal proof related to the Gumbel-Sigmoid mechanism.
  • In Section A.8, we presents the implementation details for each experiment.
  • In Section A.9, we demonstrate the generalization capability of our method with Swin Transformer liu2021swin .
  • In Section A.10, we investigate the impact of scaling up the model size to ViT-L dosovitskiy2020image .
  • In Section A.11, we provide additional visualizations of activated tokens in our appraoch.
A.1 Frequent Questions
Why the proposed method outperforms traditional adapters?
We list the explanations below:
  • The dynamic architecture in DyT enhances generalization. It introduces a form of disturbance in the input data, akin to Dropout srivastava2014dropout . This mechanism is particularly crucial when training data is limited e.g. VTAB-1K.
  • The distillation loss in DyT. We adopt the complete model as the teacher of the dynamic model, significantly enhancing performance. Such a self-distillation mechanism is only available in the dynamic architecture.
  • Previous work han2023latency and DynamicViT also show dynamic architectures outperforming static models with fewer FLOPs.
Why Table 6 shows that using MoE-adapter results in fewer FLOPs?
It is possible that MoE-adapter results in slight fewer FLOPs. We list the explanations below:
  • The FLOPs of DyT depend on learned token dispatcher during fine-tuning and may slightly fluctuate around the target FLOPs (controlled by $\mathcal{L}_{rate}$).
  • The extra computational cost of the adapters and the MoE adapters is nearly equivalent.
Thus, a DyT model with the MoE-adapter may activate fewer tokens in the learned token dispatcher, resulting in slightly reduced FLOPs.
Why the FLOPs of $N=12$ are paradoxically lower than that of $N=8$ in Table 2?
Theoretically, the MoE-adapter with any number of experts should have similar FLOPs to the standard adapter. Meanwhile, the actual activation rate of DiT during inference depends on the learned token dispatcher after fine-tuning, resulting in slight fluctuations in FLOPs between different models. These explain why DyT $N=12$ may have slight lower FLOPs than DyT $N=8$.
A.2 Difference with previous Works
We compare DyT with more previous works and demonstrate the differences between our approach and these methods.
Difference with DynamicViT and EViT.
Both DynamicViT rao2021dynamicvit and EViT liang2022not are token pruning methods, whereas DyT is a token skipping method. DynamicViT learns to retain $P$tokens at certain layers e.g. 3th,6th 9th layers in ViT-B. EViT only keeps top-K attentive tokens and fuses the inattentive tokens at certain layers e.g. 4th,7th, and 10th layer in ViT-B. These methods primarily focuses on accelerating the model within the same dataset used for pre-training, whereas DyT aims to improve efficiency during cross-domain adaptation.
Difference with DiffRate.
DiffRate chen2023diffrate is a token compression method that performs token pruning and merging simultaneously.
  • In the DiffRate chen2023diffrate , token pruning and merging is inherently data-independent. After training, a transformer layer prunes or merges the same number of tokens across all input data. In contrast, DyT is a data-dependent approach. The router in DyT learns to skip varying numbers of tokens before each MLP block based on the input data.
  • The prune and merge operations in DiffRate do not preserve complete feature maps, presenting challenges for dense prediction tasks. Thus, DiffRate requires modifications to address these tasks. Conversely, with only performing token skipping, DyT maintains complete feature maps, allowing it to effectively handle dense prediction tasks without any modifications.
Difference with ToMe.
ToMe bolya2022token is a training-free technique that enhances inference efficiency by merging tokens based on similarity at each layer. DyT employs token skipping instead of merging and can be seamlessly integrated with ToMe to further improve efficiency.
Difference with CoDA.
CoDA lei2023conditional is a PEFT method that can also improve the inference efficiency.
  • The token selection strategy in the token dispatcher is different. CoDA selects top-K tokens in each layer to pass through while DyT adopt learnable dispatchers to select an appropriate number of tokens for each input.
  • The target model is different. Although CoDA lei2023conditional also improve both the parameter and inference during adaptation, CoDA primarily focus on the language model e.g. T5 raffel2020exploring , while DyT is specifically designed for vision transformers.
  • The block to conduct token skipping is different. In CoDA, tokens directly skip the whole layer. In DyT, we propose four model variants, explore their effectiveness, and find that skipping the MLP block is the most suitable for vision transformer.
Difference with AdaMix.
AdaMix wang2022adamix also leverage a mixture of adapters, but it fuses all experts by weight averaging after training, resulting in an standard adapter akin to AdaptFormer chen2022adaptformer . In contrast, the proposed MoE-adapter employs a learning-based router to generate scalar weights for the experts based on input features, allowing features from different images to yield distinct expert weights.
A.3 More Analysis
A.3.1 Investigations of dispatch level and dispatch strategy
Token Sample TD Random CIFAR-100 SVHN Food-101 K400 SSv2
91.37 97.08 90.32 75.28 45.43
90.70 97.29 89.8 71.14 44.09
89.38 96.79 86.39 68.27 40.13
87.15 97.07 85.26 68.18 39.79
A.3.2 FLOPs-Accuracy curves of model variants.
(a) CIFAR-100
(b) SVHN
(c) Food-101
A.3.3 Different bottleneck dimensions
We investigate the impact of the bottleneck dimensions of the adapter in DyT and the results are summarized in Table 8, revealing that different datasets prefer different configurations. When the dataset is easy to adapt, such as CIFAR-100, a bottleneck dimension of $d=4$ is sufficient to achieve satisfactory performance. Conversely, video datasets require larger adapter dimensions e.g. $d=256$ to attain better performance. We set the default bottleneck dimension to 64, which achieves a balance between performance and cost, such as the number of additional parameters and FLOPs. It is worth noting that the relationship between FLOPs and bottleneck dimension is not strictly monotonic due to the existence of token dispatcher. For instance, when $d=16$, the FLOPs are lower than when $d=1$, since the training may converge at activating fewer tokens as the number of adapter’s parameters increases.
CIFAR-100 SVHN Food-101 K400 SSv2
1 0.03 12.11 91.46 95.5 89.65 71.78 36.22
4 0.08 11.89 91.57 96.61 89.99 73.16 39.24
16 0.30 12.06 91.46 96.98 90.24 73.65 42.49
64 1.19 12.21 91.37 97.08 90.32 75.28 45.43
256 4.73 13.32 91.18 97.03 90.42 75.33 46.73
A.3.4 Investigation on the Temperature
We explore the temporal $\tau$ of the proposed dynamic tuning. The results are demonstrated in Table 9. When the temperature is smaller e.g. 0.1, the Gumbel Sigmoid tends to produce binary outputs close to 0 or 1. Conversely, larger temperatures lead to more uniform outputs, approaching 0.5. Results demonstrate that the performance is not too sensitive to the temperature and our model can achieve reasonable performance with all temperatures in the table. We also observe that the model with $\tau=1$ achieves the best performance on CIFAR-100, SVHN, and SSv2, while decaying the temperature with a schedule leads to the best result on Food-101, which shows that adjusting the temperature can help the model to achieve better performance. Since identifying the optimal temperature is not the primary focus of this paper, we directly set the temperature to 5 by default.
  Temperature Image Accuracy (%) $\uparrow$ Video Accuracy (%) $\uparrow$
CIFAR-100 SVHN Food-101 K400 SSv2
0.1 90.91 96.24 89.72 73.16 44.84
1 91.61 97.20 90.08 74.38 45.69
5 91.37 97.08 90.32 75.28 45.43
Schedule 91.58 97.13 90.39 74.57 45.51
A.4 Generalization in semantic segmentation
[…]
Table 10: Results on semantic segmentation. “Avg.” denotes the average results from the corresponding two datasets. The FLOPs is measured ADE20K.
  Method Params. $\downarrow$ Semantic Segmentation Datasets
FLOPs $\downarrow$ mIOU
(M) (G) ADE20K COCO-stuff Avg.
Traditional methods
Full tuning 85.80 605.36 47.66 45.89 46.78
Linear 0 605.36 43.86 44.46 44.16
Dynamic-Full 85.80 582.46 45.74 45.20 45.47
Parameter-efficient tuning methods
AdaptFormer chen2022adaptformer 1.19 606.57 47.52 45.33 46.43
LoRA hu2021lora 1.19 605.36 47.34 45.53 46.44
VPT jia2022visual 0.07 606.36 46.38 44.87 45.63
The proposed Dynamic Tuning
DyT 1.19 584.57 46.90 45.63 46.27
[…]
A.5 Generalization in Object Detection and Instance Segmentation
We conduct experiments on COCO lin2014microsoft to explore the generalization of our method in object detection and instance segmentation. We adopt ViTDet li2022exploring as the detector and fine-tune it for 12 epochs on the COCO dataset. The bottleneck dimension $d$ in adapters is set to 128. As shown in Table 11, DyT exhibits superior performance compared to AdapterFormer chen2022adaptformer , with fewer FLOPs. Our MoE-adapter further enhances DyT without additional computational cost, validating the effectiveness of our design.
However, full tuning achieves the best performance, surpassing other methods significantly. This is likely due to the gap between bounding box regression and the pre-training of the vision transformer, necessitating more parameter updates in the backbone. This challenge motivates us to design more powerful PEFT methods and integrate them with DyT to reduce the performance gap with full tuning.
Table 11: Results on object detection and instance segmentation. We only measure the FLOPs in backbone and feature pyramid module.
  Method Params. $\downarrow$ COCO Datasets
FLOPs $\downarrow$ BBox mAP Seg mAP
(M) (G)
Full tuning 85.80 554.86 44.67 39.78
AdaptFormer chen2022adaptformer 2.56 564.53 38.71 35.27
DyT 2.56 468.40 39.78 36.11
[…]
A.6 Effectiveness of the Complete Model and Distillation
In the main paper, we adopt the complete model as our teacher, which does not employ a token dispatcher to skip tokens. Our proposed dynamic tuning allows the model to act as its own teacher during adaptation, a capability not achievable by other PEFT methods. We consider it to be a significant advantage of our approach. Joint training of the dynamic and complete models mitigates overfitting during adaptation, particularly with limited training data, such as VTAB-1K. Furthermore, the complete model, acting as a teacher, enhances the dynamic model’s learning. These factors contribute to DyT’s superior performance. Experimental results in Table 12 demonstrate that both complete model loss and distillation loss are useful for improving the performance of DyT. We also notice that introducing the complete model during training results 1.8 $\times$ longer training time. Given that our primary contribution focuses on improving parameter and inference efficiency, the additional training time introduced by the complete model would be acceptable.
In Table 13, we further present the results across all datasets in detail. We can find that, DyT without $\mathcal{L}_{cls}^{\prime}+\mathcal{L}_{distill}$ can still outperform most previous PEFT methods, further validating the superiority of our method.
Table 12: Effectiveness of loss functions The default loss function in DyT is $\mathcal{L}=\mathcal{L}_{cls}+\mathcal{L}_{cls}^{\prime}+\mathcal{L}_{distill}+\alpha\mathcal{L}_{rate}$. We gradually remove the complete model loss $\mathcal{L}_{cls}^{\prime}$ and distillation loss $\mathcal{L}_{distill}$ from it, and find the performance drops.
  Method VTAB-1K Training time
Accuracy $\uparrow$
DyT 77.14 1.8 $\times$
DyT w/o $\mathcal{L}_{distill}$ 76.70 1.8 $\times$
DyT w/o $\mathcal{L}_{cls}^{\prime}+\mathcal{L}_{distill}$ 75.73 1.0 $\times$
Table 13: Performance and efficiency comparison on VTAB-1K. “Group Mean” indicates the averaged accuracy of three groups. “Params. (M)” denotes the number of trainable parameters in backbones. “FLOPSs (G)” is the average FLOPs across all datasets. Bold font and underline denote the best and the second-best performance respectively.
Natural Specialized Structured
CIFAR-100
Caltech101
DTD
Flowers102
Pets
SVHN
Sun397
Camelyon
EuroSAT
Resisc45
Retinopathy
Clevr-Count
Clevr-Dist
DMLab
KITTI-Dist
dSpr-Loc
dSpr-Ori
sNORB-Azim
sNORB-Elev
Group Mean
Params. (M)
Traditional methods
Full tuning 68.9 87.7 64.3 97.2 86.9 87.4 38.8 79.7 95.7 84.2 73.9 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1 68.96 85.8
Frozen 63.4 85.0 63.2 97.0 86.3 36.6 51.0 78.5 87.5 68.6 74.0 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2 57.64 0
Parameter-efficient tuning methods
Adapter houlsby2019parameter 69.2 90.1 68.0 98.8 89.9 82.8 54.3 84.0 94.9 81.9 75.5 80.9 65.3 48.6 78.3 74.8 48.5 29.9 41.6 73.85 0.16
BitFit zaken2021bitfit 72.8 87.0 59.2 97.5 85.3 59.9 51.4 78.7 91.6 72.9 69.8 61.5 55.6 32.4 55.9 66.6 40.0 15.7 25.1 65.21 0.10
LoRA hu2021lora 67.1 91.4 69.4 98.8 90.4 85.3 54.0 84.9 95.3 84.4 73.6 82.9 69.2 49.8 78.5 75.7 47.1 31.0 44.0 74.60 0.29
VPT jia2022visual 78.8 90.8 65.8 98.0 88.3 78.1 49.6 81.8 96.1 83.4 68.4 68.5 60.0 46.5 72.8 73.6 47.9 32.9 37.8 71.96 0.53
SSF jie2022convolutional 69.0 92.6 75.1 99.4 91.8 90.2 52.9 87.4 95.9 87.4 75.5 75.9 62.3 53.3 80.6 77.3 54.9 29.5 37.9 75.69 0.20
NOAH zhang2022neural 69.6 92.7 70.2 99.1 90.4 86.1 53.7 84.4 95.4 83.9 75.8 82.8 68.9 49.9 81.7 81.8 48.3 32.8 44.2 75.48 0.36
ConvPass jie2022convolutional 72.3 91.2 72.2 99.2 90.9 91.3 54.9 84.2 96.1 85.3 75.6 82.3 67.9 51.3 80.0 85.9 53.1 36.4 44.4 76.56 0.33
AdaptFormer chen2022adaptformer 70.8 91.2 70.5 99.1 90.9 86.6 54.8 83.0 95.8 84.4 76.3 81.9 64.3 49.3 80.3 76.3 45.7 31.7 41.1 74.75 0.16
FacT-TT jie2023fact 71.3 89.6 70.7 98.9 91.0 87.8 54.6 85.2 95.5 83.4 75.7 82.0 69.0 49.8 80.0 79.2 48.4 34.2 41.4 75.30 0.04
Res-Tuning jiang2023res 75.2 92.7 71.9 99.3 91.9 86.7 58.5 86.7 95.6 85.0 74.6 80.2 63.6 50.6 80.2 85.4 55.7 31.9 42.0 76.32 0.51
The proposed Dynamic Tuning without $\mathcal{L}_{cls}^{\prime}+\mathcal{L}_{distill}$
DyT $r=0.5$ 70.4 94.2 71.1 99.1 91.7 88.0 51.5 87.1 95.3 84.2 75.8 79.2 61.8 51.0 82.4 79.7 52.3 35.3 44.5 75.73 0.16
DyT $r=0.7$ 73.9 94.9 72.1 99.4 91.8 88.4 55.5 87.2 95.6 86.2 75.9 80.3 61.8 51.7 83.1 81.6 53.7 35.3 45.2 76.69 0.16
DyT $r=0.9$ 74.0 95.1 72.9 99.3 91.7 87.6 56.9 87.7 95.7 85.4 76.1 81.6 63.2 50.1 83.0 83.3 52.0 34.5 44.5 76.74 0.16
The proposed Dynamic Tuning with distillation
DyT $r=0.5$ 73.6 94.8 73.0 99.1 91.4 87.0 56.4 87.3 96.1 85.6 76.7 82.8 63.8 52.7 83.7 83.6 57.3 34.6 44.3 77.14 0.16
DyT $r=0.7$ 74.4 95.5 73.6 99.2 91.7 87.5 57.4 88.3 96.1 86.7 76.7 83.5 63.8 52.9 83.1 85.7 54.9 34.3 45.9 77.57 0.16
DyT $r=0.9$ 74.3 94.9 73.1 99.2 91.4 87.8 57.1 87.9 96.1 85.9 76.0 83.3 64.8 51.5 83.4 84.0 54.8 35.1 46.4 77.30 0.16
A.7 Details about Gumbel-Sigmoid